The bond length of HCl molecules is 1.275A∘1.275A∘ and its dipole moment is 1.03D1.03D. The ionic character of the molecule (in percent) (charge of the electron =4.8×10−10e.s.u)=4.8×10−10e.s.u) is(1)100(3)33.66(2)67.3(4)16.83

Grade:12th pass

1 Answers

Aastha Singh
5 years ago

Hello student,

convert all the given data into same units

bond length = 1.275 * 10-8 c.m.
charge on electron = 4.8 * 10-10 esu
ideal dipole moment = 1.275 * 10-8 * 4.8 * 10-10
=
6.12 * 10-18 esu cm
observed dipole moment = 1.03 Debye = 1.03 * 10-18 esu c.m.
% ionic character = observed dipole moment /ideal dipole moment * 100
1.03 *10-18/6.12 * 10-18 = 0.1683* 100
=16.83